# Nonuniform Sobolev Spaces

### 05 Jan 2024 v1

@@ 4. Convergence of Schrödinger Operators
Note that $(f*\varphi)\,\hat{}=\hat{\varphi}\hat{f}$ (for a proof, see [28, Theorem 7.19]). Hence $f_{1}=f*\varphi$ and $f_{2}=f-f*\varphi$. Now we rewrite $e^{it(-\Delta)^{a/2}}$ as
@@
    If $20<s<N/$ 2, then (4.2) is true for all functions $f\in W_{s}^{\vec{p}}$.
@@
    $e^{-it\Delta}f(x)=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{i(x\cdot\omega+t\omega^{2})}\Big{(}1-\varphi_{0}(\frac{\omega}{\varepsilon})\Big{)}\hat{f}(\omega)\mathrm{d}\omega,$ (4.3)
@@
If $10<s<$ 1, $p_{0}>N/\alpha>p_{1}=2$ and $2(\alpha+s)>N$, then the function $f$ defined in Example 3.6 satisfies that $f\in W_{s}^{(p_{0},2)}\setminus W_{s}^{(2,2)}$. That is, $W_{s}^{(p_{0},2)}$ is not contained in $W_{s}^{(2,2)}=H^{s}$.
@@
There are three cases.
@@
$e^{-it\Delta}f(x)$ $=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{f}(\omega)\mathrm{d}\omega.$ (4.22)
@@
$\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{f}(\omega)\mathrm{d}\omega$
$=\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{\varphi}(\omega)\hat{f}(\omega)\mathrm{d}\omega+\int_{|\omega|>1}e^{i(x\omega+t\omega^{2})}\hat{f}_{2}(\omega)\mathrm{d}\omega.$
@@
$e^{-it\Delta}f_{1}(x)$ $=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{\varphi}(\omega)\hat{f}(\omega)\mathrm{d}\omega$
$=\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{i(x\omega+t\omega^{2})}\hat{\varphi}(\omega)\rangle.$
@@
Observe that $e^{i(x\omega+t\omega^{2})}\hat{\varphi}(\omega)$ tends to $e^{ix\omega}\hat{\varphi}(\omega)$ in $\mathscr{S}$ as $t\rightarrow 0$. We get
$\lim_{t\rightarrow 0}e^{-it\Delta}f_{1}(x)$ $=\lim_{t\rightarrow 0}\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{i(x\omega+t\omega^{2})}\hat{\varphi}(\omega)\rangle$
$=\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{ix\omega}\hat{\varphi}(\omega)\rangle$

### 23 Jan 2024 v3

@@ 4.2. Convergence of Schrödinger Operators
Note that $(f*\varphi)\,\hat{}=\hat{\varphi}\hat{f}$ (for a proof, see [37, Theorem 7.19]). Hence $f_{1}=f*\varphi$ and $f_{2}=f-f*\varphi$. Now we rewrite $e^{it(-\Delta)^{a/2}}$ as
@@
    If $20<s<N/$ 2, then (4.39) is valid for all functions $f\in W_{s}^{\vec{p}}$.
@@
    $e^{-it\Delta}f(x)=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{i(x\cdot\omega+t|\omega|^{2})}\Big{(}1-\varphi_{0}(\frac{\omega}{\varepsilon})\Big{)}\hat{f}(\omega)\mathrm{d}\omega,$ (4.40)
@@
If $10<s<$ 1, $p_{0}>N/\delta>p_{1}=2$ and $2(\delta+s)>N$, then the function $f$ defined in Example 3.6 satisfies that $f\in W_{s}^{(p_{0},2)}\setminus W_{s}^{(2,2)}$. That is, $W_{s}^{(p_{0},2)}$ is not contained in $W_{s}^{(2,2)}=H^{s}$.
@@
There are three cases:
@@
$e^{-it\Delta}f(x)$ $=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{|\omega|>\varepsilon}e^{i(x\cdot\omega+t|\omega|^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{f}(\omega)\mathrm{d}\omega.$ (4.59)
@@
$\int_{|\omega|>\varepsilon}e^{i(x\cdot\omega+t|\omega|^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{f}(\omega)\mathrm{d}\omega$
$=\int_{|\omega|>\varepsilon}e^{i(x\cdot\omega+t|\omega|^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{\varphi}(\omega)\hat{f}(\omega)\mathrm{d}\omega+\int_{|\omega|>1}e^{i(x\cdot\omega+t|\omega|^{2})}\hat{f}_{2}(\omega)\mathrm{d}\omega.$
@@
$e^{-it\Delta}f_{1}(x)$ $=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{|\omega|>\varepsilon}e^{i(x\cdot\omega+t|\omega|^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{\varphi}(\omega)\hat{f}(\omega)\mathrm{d}\omega$
$=\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{i(x\cdot\omega+t|\omega|^{2})}\hat{\varphi}(\omega)\rangle.$
@@
Observe that $e^{i(x\cdot\omega+t|\omega|^{2})}\hat{\varphi}(\omega)$ tends to $e^{ix\cdot\omega}\hat{\varphi}(\omega)$ in $\mathscr{S}$ as $t\rightarrow 0$. We get
$\lim_{t\rightarrow 0}e^{-it\Delta}f_{1}(x)$ $=\lim_{t\rightarrow 0}\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{i(x\cdot\omega+t|\omega|^{2})}\hat{\varphi}(\omega)\rangle$
$=\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{ix\cdot\omega}\hat{\varphi}(\omega)\rangle$